# VersiCode: Towards Version-controllable Code Generation

### v1

@@ Abstract
Significant research has focused on improving the performance of large language model on code-related tasks due to their practical importance. Although performance is typically evaluated using public benchmark datasets, the existing datasets do not account for the concept of version, which is crucial in professional software development. In this paper, we introduce VersiCode, the first comprehensive dataset designed to assess the ability of large language models to generate verifiable code for specific library versions. VersiCode encompasses 300 libraries across more than 2,000 versions spanning 9 years. We design two dedicated evaluation tasks: version-specific code completion (VSCC) and version-aware code editing (VACE). Comprehensive experiments are conducted to benchmark the performance of LLMs, revealing the challenging nature of these tasks and VersiCode, that even state-of-the-art LLMs struggle to generate version-correct code. This dataset, together with the proposed tasks, sheds light on LLMs’ capabilities and limitations in handling version-specific code generation, and opens up an important new area of research for further investigation. The resources can be found at https://github.com/wutong8023/VersiCode.
@@ 1 Introduction
Large language models (LLMs) [35] and their code-specific variants, such as CodeLlaMa [37], have shown strong capabilities in code generation tasks. Many code generation datasets have been proposed [6, 27] to evaluate these models’ performance, which sees steady increases. While these datasets serve as important benchmarks, they all treat code generation as a static task [18], without considering the influence of versioning of dependent libraries. In contrast, in professional software development, as shown in Figure 1, dependency management is an essential practice that plays a key role in building dependable software. Keeping track of versions is a central step in dependency management. Given the evolution of software libraries and their APIs, two major challenges remain under-explored. 1) How to generate code that uses and fits a specific version of a library. 2) How to adapt code using an older version of a library to a newer version.
@@
In this paper, we propose two dedicated version-controllable code generation tasks to simulate realistic settings in professional software development: version-specific code completion (VSCC) and version-aware code editing (VACE), in which the versions of the libraries serve as a crucial constraint for code generation.
Existing code generation datasets, due to their obliviousness to version information, cannot support these two tasks [6, 27]. Thus, to evaluate LLMs’ performance on these two novel tasks, we curate and release VersiCode, a large-scale benchmark dataset in Python with comprehensive versioning metadata. VersiCode contains 300 libraries and 2,207 versions, spanning 9 years from 2015 to 2023. We carefully developed a human-LLM hybrid data collection and annotation pipeline to create a large and high-quality dataset. Each data instance in VersiCode is represented as a tuple of “<library version, functionality description, code snippets>”. Subsequently, by detecting version-sensitive parts (e.g. function or variable names), we post-processed the metadata into tasks of varying granularities: token-level, line-level, and block-level, reflecting different levels of difficulty by generating more content and reducing provided contextual cues. To the best of our knowledge, VersiCode is the first version-controllable benchmark dataset. Our new benchmark dataset VersiCode, viewed through the lens of library version evolution, addresses the following research questions:
  1. 1.
    How well do existing LLMs understand version requirements for code generation?
  2. 2.
    Can LLMs recall the version span of APIs in each library, i.e., from addition to deprecation? Furthermore, can LLMs reason across intermediate versions?
  3. 3.
    What is the trend of LLM performance regarding the release time of each library?
In our comprehensive experiments, we evaluated the performance of a large number of state-of-the-art LLMs, including open LLMs such as Llama 2 and Llama 3, proprietary LLMs such as GPT-3.5 and GPT-4o, as well as seven code-specific variants such as CodeLlama-13B, StarCoder2-15B, and CodeQwen1.5-7B. On the simplest token-level code completion task, these LLMs perform significantly worse on VersiCode than on other well-known datasets including HumanEval(+) and MBPP(+). For some LLMs, the performance gap of Pass@1 between existing benchmarks and VersiCode is more than 50 points. Even the most capable LLM, GPT-4o, obtains a Pass@1 score of 70.44 on VersiCode while achieving a score higher than 85 on HumanEval and MBPP. LLMs performance on line-level and block-level completion is in generally much lower than that on token-level. For the version-aware code editing task, there is significant room for improvement even for the most capable LLM GPT-4o. Our contributions are summarized as follows:
@@
  • We propose two novel and important yet under-explored tasks in code generation, namely version-specific code completion (VSCC) and version-aware code editing (VACE).
  • We curate VersiCode, a new, large-scale version-controllable code generation dataset, aimed at comprehensively evaluating LLMs’ performance on the above two tasks.
  • We conduct comprehensive experiments on VersiCode, demonstrating that it is a high-quality and challenging dataset. The benchmarking results reveal that most LLMs struggle with version-specific code generation, especially with the latest libraries, highlighting areas for future exploration.
@@
2 VersiCode: A Benchmark for Version-controllable Code Generation
VersiCode is a large-scale code generation benchmark dataset focusing on evolving library dependencies. We propose two tasks to simulate real-world applications: version-specific code completion and version-aware code editing, incorporating version information into code generation constraints. First, we discuss data curation, preprocessing of noisy code snippets and FAQs into organized metadata (Section 2.1). Based on this metadata, we describe the task design and quality control process (Section 2.2). We then address tagging API lifespan features per library version (Section 2.3). Finally, we provide data statistics for VersiCode and discuss future dataset extensions (Section 2.4).
2.1 Dataset Curation and Collection
As shown in Figure 2, we first collected permissively licensed Python repositories from GitHub that serve as the source code for Python libraries. These repositories are ranked by their popularity (as indicated by their collected stars). Using the list of popular libraries, we gathered data from three sources for each library: (1) Library Source Code: We collected all available versions of the library source code from GitHub, verifying with PyPI to ensure that the collected versions are formally released and can be installed via pip. From the library source code, we extracted official usage examples for each API from the docstrings. (2) Downstream Application Code: Given Python’s popularity in scientific programming, we collected the source code from top-tier research papers over 10 years as downstream applications. These applications are valuable due to being lightweight yet self-consistent, diverse in their topics, and tagged release timelines associated with publishing venues. Given the time span, this data source implicitly includes evolving libraries. (3) Stack Overflow: Using the library names as queries, we collected FAQ data from Stack Overflow, which provides real user queries and diverse user answers. We filtered the data to include only those queries that explicitly mention the versions of the libraries used, using heuristic rules, refer to Appendix A.
Figure 2: The preprocessing pipeline to obtain metadata, structured as n-gram tuple of $<$library name, version, functionality description, code snippet$>$.
2.2 Task Design for Version-controllable Code Generation
As shown in Figure 3, we define each meta-instance $m_{i}=[l_{i};v_{i};d_{i};c_{i}]\in\mathcal{M}$, where $l$, $v$, $d$, and $c$ represent the library name, version, functionality description, and code snippet, respectively. We then design the following two version-controllable code generation tasks:
@@ 2.2 Task Design for Version-controllable Code Generation
Version-Aware Code Editing (VACE): Given a pair of meta-instances $(m_{i},m_{j}|l_{i}==l_{j},d_{i}==d_{j},v_{i}\neq v_{j})$, the input $x=[l_{i};v_{i};d_{i};c_{i};v_{j}]$, and the output $y=c_{j}$. Note that version editing may require refactoring of the code structure, making it difficult to format as detailed as in token-level or line-level completion. Additionally, depending on the numerical relationship between $v_{i}$ and $v_{j}$, various scenarios arise, such as editing from an old version to a new version, or vice versa.
2.3 Lifespan Tagging of APIs
@@
Figure 3: The post-processing pipeline transforms metadata into specific tasks and the running example per task: (left) Leveraging pairs of metadata that share the same functionality but different library versions to construct block-level code editing instances; (right) Utilizing each metadata sample, masking version-sensitive content to create multi-granularity code completion instances.
2.4 Dataset Statistics, Scope, and Future Extensions
Defects4J [20] Open Source Programs Java 357 Debug Block-Level 2014 None
QuixBugs [26] Quixey Challenges Python, Java 40 Debug Line-Level 2017 Human
CoST [55] GeeksForGeeks Multi(=7) 132,046 Code Translation Line-Level, Block-Level 2022 None
XLCoST [54] GeeksForGeeks Multi(=8) 1,083,000 Code Translation Line-Level, Block-Level 2022 None
InstructCoder [15] Github Python 114,000 Code Refinement Block-Level 2023 LLM
MultilingualTrans [47] Programming Sites Multi(=8) 30,419 Code Translation Block-Level 2023 None
NicheTrans [47] Programming Sites Multi(>8) 236,468 Code Translation Block-Level 2023 None
LLMTrans [47] Hand-Written Multi(=8) 350 Code Translation Block-Level 2023 Human
Avatar [2] Programming Sites Python, Java 62,520 Code Translation Block-Level 2023 None
G-TransEval [19] Existing benchmark, GeeksForGeeks Multi(=5) 400 Code Translation Token-Level, Block-Level 2023 Human
EvalGPTFix [51] AtCoder Java 151 Debug Block-Level 2023 Human
DebugBench [41] LeetCode Multi(=3) 4,253 Debug Block-Level 2024 LLM
VersiCode Github Python 98,692 Version Adaptation Block-Level 2024 LLM
Dataset Statistics: We present the statistics of VersiCode in Table 2, using the NLTK  [^1] https://www.nltk.org/ tokenizer to compute the number of tokens. As shown in Table 1, we also provide a comparison between VersiCode and other code editing datasets. See Appendix B.2 for the comparison between VersiCode and other code completion datasets.
Scope: VersiCode supports version-specific code completion at the token, line, and block levels, enabling developers to navigate through version variations effortlessly. It also facilitates block-level version-aware code editing, empowering users to make precise modifications tailored to requirements of each version. The collected metadata also serves as a valuable resource for potential customized task modifications, aiding in fine-tuning workflows and enhancing model training for optimal performance.
Table 2: Data statistics of VersiCode.
# Num. of Libraries 300
# Num. of Versions 2,207
# Size of Metadata 11,268
# Task Type Completion Editing (old to new) Editing (new to old)
# Granularity Token Line Block Block Block
# Avg. Input Token 1,233 1,213 44 115 116
# Avg. Output Token 1 9 77 70 69
# Num. of Instances 13,532 13,530 1,618 49,346 49,346
Future Extension: VersiCode primarily supports Python, but we have also made efforts to extend it to multiple languages. Refer to the Appendix B.1 for comprehensive data statistics and initial multi-language evaluation results. Given the rapid evolution of software technologies, continued effort to further automate the curation pipeline and maintenance of VersiCode is needed.
3 Experiments
We undertake experiments to investigate the following research questions:
  1. 1.
    What is the quality and difficulty of VersiCode, particularly in comparison to related datasets and among diverse data sources?
  2. 2.
    What is the current proficiency of codeLLMs in generating version-controllable code, considering comparisons among multiple-granularity completions, for both code completion and code editing?
  3. 3.
    What is the timeliness of LLMs’ programming knowledge in code generation, with a focus on model performance categorized by year over the past decade?
3.1 Implementation Details
@@
Models: We benchmarked VersiCode against popular open-domain LLMs and dedicated code-LLMs, including variant families such as GPT [35, 34, 36], LLaMa [42], Mistral [17], CodeLLaMa [37], CodeQwen [5], CodeGemma [7], StarCoder [28], Deepseek-Coder [13], and WizardCoder [32]. For smaller open-source models (e.g., < 20B parameters), we downloaded them from HuggingFace  [^2] https://huggingface.co/models and deployed them locally for inference. For larger models, such as LLaMa3 70B [33] and GPT-4o [36], we used their online APIs  [^3] https://together.ai  [^4] https://openai.com/ for inference.
@@
Baseline Dataset: To assess the difficulty of VersiCode, we compared it with two well-known code generation datasets, HumanEval [27] and MBPP [18], and observed the overall performance of models. HumanEval [27] measures functional correctness in synthesizing programs from docstrings with 164 original problems, resembling simple software interview questions. MBPP [4], with about 1,000 crowd-sourced Python problems for entry-level programmers, covers programming fundamentals and standard library functionality, including task descriptions, code solutions, and three automated test cases for each problem. We also collected the evaluation results for their upgraded versions HumanEval+ [27] and MBPP+ [27]. Please refer to Appendix E.1 for details.
Data Preparation: As introduced in Section 2.2, we designed two types of version-controllable code generation tasks: version-specific code completion and version-aware code editing. The task granularities are categorized into token-level, line-level, and block-level to control difficulty and simulate different application scenarios. To better understand model performance, each instance in VersiCode is also tagged with the following: (1) data source, which includes library source code, downstream applications, and Stack Overflow (see Section 2.1); (2) feature type, including addition, deprecation, and general (see Section 2.3); (3) release time, i.e. the timestamp from GitHub and Stack Overflow); and (4) considering code editing instances constructed from pairs of metadata, the differences between source and target code versions result in various situations, such as updates from an older version to a newer version or vice versa. Additionally, we categorized versions according to version patterns, for example, treating torch v1.0.0 as a major version and torch v1.3.1 as a minor version, to identify combinations of major and minor version editing cases. These tags allow us to filter the evaluation dataset and gain sharper insights into model performance.
@@
Evaluation Metrics: We use the following evaluation metrics for each task granularity: (1) Pass@$k$ for token-level generation [6]: For this metric, we generate $n\geq k$ samples per instance (with $n=100$ and $k\in\{1,3,10\}$ for our experiments). We count the number of correct samples $c\leq n$ judged by exact matching. Pass@$k$ is defined as the average performance over the task, calculated as $\mathbb{E}\left[1-\frac{{\binom{n-c}{k}}}{\binom{n}{k}}\right]$. (2) Identifier Sequence Match (ISM@$n$) and Prefix Match (PM@$n$) for line-level generation [1]: These metrics measure how closely the generated sequences match the ground truth. For block-level generation, we adopt the average performance over lines. Following the setup in Agrawal et al. [1], we generate $n=6$ independent samples per instance.
@@ 3.2 Results and Analysis
Figure 4: Pass@1 results for token-level code completion from VersiCode: (a1) comparison with existing benchmark datasets, (a2) performance grouped by data sources, and (b) performance grouped by API lifespan features.
@@
Even token-level code completion is challenging. We present the pass@1 results of token-level code completion for LLMs on VersiCode, sorted by release time (see Figure 4-a1, highlighted in green). When compared to the Pass@1 results on HumanEval (in blue) and MBPP (in orange), all models perform significantly worse on VersiCode (in green). This result indicates the difficulty in disambiguating and recalling version-specific library usage. It is worth noting that the larger and latest models, such as GPT-4o (M13) and LLaMA3-70B (M12), achieve significantly better performance than the other models (See Appendix D.1 for the error analysis of GPT-4o.). However, the performance gap with HumanEval and MBPP is still large, with at least 15 points. Thus, for the simplest token-level completion task, state-of-the-art LLMs struggle to achieve satisfactory performance.
@@
Differences in LLM performance across different data sources. We present the Pass@1 results, categorized by data sources, of token-level code completion for LLMs on VersiCode in Figure 4-a2. Comparing these three data sources, most models perform much better on Stack Overflow than on the other two, especially the source code from downstream applications. This result may be due to the high diversity in downstream applications, which requires a strong ability to tackle them. It may also suggest that Stack Overflow is highly represented in the pre-training data of LLMs, hence a greater chance of data leakage. Similar to Figure 4-a1, the outliers are still GPT-4o (M13) and LLaMA3-70B (M12), which excel in dealing with downstream applications increasing the likelihood of models memorizing specific content. Please refer to Appendix E.1 for full numeric results.
@@
Challenges in casual intermediate library versions. We present the token-level Pass@1 results categorized by lifespan features: addition (in blue), deprecation (in orange), and general (referring to intermediate versions; in green) for the token-level code completion task (see Figure 4-b). Most models perform well in the cases of addition and deprecation due to newly added or deprecated APIs, as these versions are likely emphasized in documentation or by the community. However, most models struggle with reasoning and adapting to intermediate versions. When viewed alongside Figure 4-a2, it is evident that models such as LLaMA3-70B perform better in downstream applications and are also good at intermediate versions, benefiting from the diversity of use cases.
Table 3: Evaluation of multi-granularity version-specific code completion, and version-adapted code editing in both “old $\mapsto$ new” and “new $\mapsto$ old” scenarios.
Model Code Completion Code Editing (old $\mapsto$ new) Code Editing (new $\mapsto$ old)
Token-level Line-level Block-level Block-level Block-level
Pass@1 Pass@3 Pass@10 ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6
DeepSeek-Coder-7B-Instruct-V1.5 [13] 15.71 21.72 27.84 35.59 22.53 0 0 45.41 36.44 40.25 28.89
CodeLlama-13B-Instruct [37] 13.5 20.29 26.87 29.22 16.69 4.34 3.12 39.17 27.94 39.12 26.17
StarCoder2-15B [28] 14.55 21.35 27.59 20.37 12.46 1.36 0.79 35.59 26.72 41.41 27.64
CodeGemma-7B [7] 13.28 18.79 24.09 5.2 2.57 0.37 0.22 9.16 4.12 9.72 5.28
GPT-3.5-Turbo [34] 37.59 44.62 51.17 49.19 32.85 40.28 27.57 45.96 36.8 46.96 35.06
Llama-3-70B-Chat [33] 47.55 52.42 55.77 49.62 33.14 57.68 41.47 33.37 23.51 42.94 29.36
GPT-4o [36] 70.44 74.27 77.27 69.44 58.95 64.73 50.48 54.72 45.04 55.36 52.33
Reduced context increases error risk in code generation. Based on the token-level code completion performance of each model, we selected the top models for further analysis. We present the multi-granularity comparison in Table 3. When comparing the performance between line-level and block-level code completion, we can observe that smaller models tend to fail more frequently at generating correct code in block-level completion, due to less code context and the requirement to generate more content, which aligns with our intuition. Note that the results shown here have been filtered by grammar verification, a post-generation validation step that only counts code that successfully compiles in Python. If we remove grammar verification, the overall performance of block-level completion in Table 11 (Appendix E.3) is comparable to line-level completion in Table 3. This suggests that while the models can predict code-style content, they cannot guarantee correct programming grammar.
The context code in another version is still helpful, but its benefits are limited. The comparison between block-level code completion and block-level code editing is shown in Table 3. There is a significant improvement across most models, except for LLaMA3-70B and GPT4-o. When provided with code in another version as context (i.e. in the code editing task), these models can generate correct code with a much higher success rate. However, a bottleneck is evident in LLaMA3-70B and GPT4-o, where the code context hinders their performance compared to code completion.
Major version matters in version-aware code editing. As shown in Table 3, “old to new” editing and “new to old” demonstrate similar performance among models. As shown in Figure 4, we categorize editing instances according to their source and target versions, distinguishing between major and minor versions. It’s evident that when the major version serves as the source, the model’s editing performance is inferior compared to other scenarios.
Table 4: Performance of the model’s editing capabilities, influenced by transitions between source and target versions, such as shifts from major to minor versions or vice versa.
[…]
ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6 ISM@6 PM@6
DeepSeek-Coder-7B-Instruct-V1.5 [13] 12.83 11.10 23.73 18.91 30.71 22.97 36.29 21.82
CodeLlama-13B-Instruct [37] 15.4 14.2 23.07 18.29 29.10 20.65 34.09 20.36
StarCoder2-15B [28] 14.94 13.40 19.90 14.87 24.74 16.99 35.10 21.91
CodeGemma-7B [7] 7.71 5.89 7.10 6.03 6.00 3.54 13.45 7.93
GPT-3.5-Turbo [34] 14.47 12.40 24.57 20.60 32.07 25.20 39.57 25.57
Llama-3-70B-Chat [33] 15.50 14.10 20.47 17.00 21.46 16.35 32.40 19.77
GPT-4o [36] 26.21 24.60 41.39 37.23 50.01 44.20 59.55 47.40
@@
Figure 5: Pass@1 performance for token-level code completion, grouped by year (2015-2023), with a histogram of data distribution for each year.

### v2

# Tongtong Wu     Weigang Wu      Xingyu Wang      Kang Xu     Suyu Ma     Bo Jiang Ping Yang     Zhenchang Xing     Yuan-Fang Li     Gholamreza Haffari

@@ Abstract
Large Language Models (LLMs) have made tremendous strides in code generation, but existing research fails to account for the dynamic nature of software development, marked by frequent library updates. This gap significantly limits LLMs’ deployment in realistic settings. In this paper, we propose two novel tasks aimed at bridging this gap: version-specific code completion (VSCC) and version-aware code migration (VACM). In conjunction, we introduce VersiCode, a comprehensive Python dataset specifically designed to evaluate LLMs on these two tasks, together with a novel evaluation metric, Critical Diff Check (CDC@1), which assesses code generation against evolving API requirements. We conduct an extensive evaluation on VersiCode, which reveals that version-controllable code generation is indeed a significant challenge, even for GPT-4o and other strong frontier models. We believe the novel tasks, dataset and metric open up a new, important research direction that will further enhance LLMs’ real-world applicability. The code and resources can be found at https://wutong8023.site/VersiCode/.
@@ 1 Introduction
Large Language Models (LLMs), including OpenAI’s GPT series (OpenAI, 2023a; b; 2024) and specialized variants such as CodeLLaMA (Rozière et al., 2023), have demonstrated significant advancements in code generation tasks. Typically evaluated using benchmarks such as HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021), these models are measured on tasks that assume code generation is a static activity. However, the reality of software development is inherently dynamic, characterized by frequent updates to software libraries, which necessitate adjustments to API interfaces. This evolving landscape raises crucial challenges for LLMs, particularly their ability to generate code that is functional for different, specific library versions. This dynamic nature of software development leads us to ask the following questions:
  • How reliably can LLMs generate code compatible with specific library versions?
  • How effectively can LLMs adapt code for API changes across library versions?
Existing benchmarks (Jiang et al., 2024; Sun et al., 2024; Luo et al., 2024b), which are oblivious to version-specific dynamics, do not address these challenges. They fall short of simulating the continuous version management activities undertaken by developers who ensure the software remains functional across updates. The static nature of existing benchmarks represents a significant barrier to the practical deployment of LLMs in professional environments, where handling version-specific dependencies is critical (Zhang et al., 2020; 2021; Dilhara et al., 2021; Liu et al., 2021; Wang et al., 2020; Vadlamani et al., 2021; Haryono et al., 2021).
To bridge this gap, we propose two novel tasks aimed at evaluating LLMs’ version-controllable code generation capabilities, namely version-specific code completion (VSCC) and version-aware code migration (VACM). These tasks are crafted to mimic real-world software development scenarios, motivated in Figure 1, requiring models to generate code that not only is syntactically correct but also adheres to version-specific API contracts (Zhang et al., 2020; 2021; Dilhara et al., 2021; Liu et al., 2021; Wang et al., 2020; Vadlamani et al., 2021; Haryono et al., 2021). Moreover, we introduce VersiCode, the first dataset specifically designed for these two tasks. VersiCode includes data spanning over 300 Python libraries and more than 2,000 versions across 9 years. It has undergone a careful curation process to ensure high quality. Thus, VersiCode provides a comprehensive and robust testbed for assessing LLMs under realistic conditions. Furthermore, we propose a new evaluation metric, CDC (Critical Diff Check), which enhances traditional code similarity metrics by incorporating considerations for API usage, parameter handling, and deprecated features management. This metric offers a more granular assessment of a model’s ability to navigate the complexities of evolving software libraries.
Our extensive testing of strong frontier models like GPT-4o and LLaMA3 (Meta LlaMa team, 2024) reveals significant challenges in version-aware code generation tasks. We uncover that (1) LLMs often retain outdated programming knowledge, particularly concerning version-specific information. (2) Conventional metrics used for evaluating code generation do not effectively capture the nuances of version sensitivity. (3) While leveraging context from various library versions can be beneficial, its utility can be limited. Guided by these insights, we suggest strategies, such as targeted pretraining, continual learning, and refined evaluation methods, for improving LLMs’ version-controlled code generation capabilities.
@@
Our contributions are summarized as follows:
@@
  • We propose two novel and important yet under-explored tasks in code generation, namely version-specific code completion and version-aware code migration.
  • We introduce VersiCode, a comprehensive, well-documented and versioned dataset, accompanied by a subset annotated with executable test cases.
  • We introduce Critical Diff Check, a new metric that extends traditional code similarity metrics by checking syntactic validity, API usage, parameter matching, the use of ‘with‘ statements, and correct keyword arguments in the generated code, providing a more detailed evaluation of version-specific code generation.
  • Our thorough experiments provide valuable insights and directions for future research in this critical area of software development.
2 Version-controllable Code Generation
VersiCode is a large-scale code generation benchmark dataset focusing on evolving library dependen- cies. We curated our dataset by initially selecting popular Python repositories from GitHub, confirmed by their star ratings, and ensured they were permissively licensed. For each library, we compiled data from three main sources: (1) Library Source Code, extracting all pip-installable versions and official API usage examples from docstrings; (2) Downstream Application Code, sourcing from top-tier research papers spanning ten years to capture evolving libraries; (3) Stack Overflow, retrieving FAQs that mention specific library versions. We present the dataset statistics, construction process and examples in detail in Appendix 2.
As shown in Figure 2, we define a meta-instance as $m=[l;v;d;c]\in\mathcal{M}$, where $l$, $v$, $d$, and $c$ represent the library name, version, functionality description, and code snippet, respectively. Consider an API $a$ added to library $l$ in version $v_{s}$ and deprecated in version $v_{e}$, and is active in the intermediate version $v_{m}$ where $s\leq m\leq e$. We refer to the interval $[s,e)$ as the lifecycle of $a$. To analyze model performance in detail, we assess how up-to-date each LLM is concerning newly added or deprecated APIs per version. We compare the source code between any two consecutive versions of each library to detect changes in API or method names. Based on the detection results, we label the source code as follows: “addition” indicates an API newly added in the current version and still applicable in subsequent versions; “deprecation” indicates the current version is the last usable version for the API; and “general” indicates the API usage method is inherited from the previous version.
@@ 2 Version-controllable Code Generation
We introduce the two novel version-controllable code generation tasks below.
@@
Version-Aware Code Migration (VACM): Given a pair of meta-instances $(m_{i},m_{j}|l_{i}==l_{j},d_{i}==d_{j},v_{i}\neq v_{j})$, the input $x=[l_{i};v_{i};d_{i};c_{i};v_{j}]$, and the output $y=c_{j}$. Note that version editing may require refactoring of the code structure, making it difficult to format as detailed as in token-level or line-level completion. Additionally, depending on the numerical relationship between $v_{i}$ and $v_{j}$, various scenarios arise, such as editing from an old version to a new version, or vice versa. Data statistics are detailed in Appendix B
@@
Figure 2: The post-processing pipeline transforms metadata into specific tasks and the running example per task: (left) Leveraging pairs of metadata that share the same functionality but different library versions to construct block-level code migration instances; (right) Utilizing each metadata sample, masking version-sensitive content to create multi-granularity code completion instances.
3 Token-level Version-specific Code Completion
In code generation that targets a specific version of a third-party library, the version-related changes usually involve updates to identifiers, such as the addition, removal, or renaming of classes, functions, and parameters. The token-level code completion task for a specified version, predicting the evolving identifiers identified in real code, is a fundamental and direct way to evaluate LLMs to generate code for specific versions. We begin our research by addressing the following three research questions: (1) How well do LLMs perform on code completion tasks that involve version-specific library usage compared to other benchmarks like HumanEval and MBPP? (2) How do LLMs handle new, deprecated, and intermediate versions of libraries in code completion tasks? (3) How does the performance of LLMs in code completion change over time with different library versions?
3.1 Experiment Setup
@@
Models: We benchmarked VersiCode against popular open-domain LLMs and dedicated code-LLMs, including variant families such as GPT (OpenAI, 2023a; b; 2024), LLaMa (Touvron et al., 2023), Mistral (Jiang et al., 2023), CodeLLaMa (Rozière et al., 2023), CodeQwen (Bai et al., 2023), CodeGemma (CodeGemma Team et al., 2024), StarCoder (Lozhkov et al., 2024), Deepseek-Coder (Guo et al., 2024), and WizardCoder (Luo et al., 2024c). For smaller open-source models (e.g., <20B parameters), we downloaded them from HuggingFace  [^1] https://huggingface.co/models and deployed them locally for inference. For larger models, such as LLaMa3 70B (Meta LlaMa team, 2024) and GPT-4o (OpenAI, 2024), we used their online APIs  [^2] https://together.ai  [^3] https://openai.com/ for inference.
Data Preparation: Each instance in VersiCode is tagged with its data source (library source code, downstream applications, or Stack Overflow), feature type (addition, deprecation, or general), and release time, allowing for more detailed performance analysis. We randomly selected 2,000 instances for token-level code completion. (see Appendix A.3).
@@
Baseline Dataset: To assess the difficulty of VersiCode, we compared it with two well-known code generation datasets, HumanEval (Liu et al., 2023) and MBPP (Jiang et al., 2024), and observed the overall performance of models. HumanEval (Liu et al., 2023) measures functional correctness in synthesizing programs from docstrings with 164 original problems, resembling simple software interview questions. MBPP (Austin et al., 2021), with about 1,000 crowd-sourced Python problems for entry-level programmers, covers programming fundamentals and standard library functionality, including task descriptions, code solutions, and three automated test cases for each problem. We also collected the evaluation results for their upgraded versions HumanEval+ (Liu et al., 2023) and MBPP+ (Liu et al., 2023). Please refer to Appendix D.1 for details.
@@
Evaluation Metrics: We use EM@$k$ for token-level generation: For this metric, we generate $n\geq k$ samples per instance (with $n=100$ and $k\in\{1,3,10\}$ for our experiments). We count the number of correct samples $c\leq n$ judged by exact matching. @$k$ is defined as the average performance over the task, calculated as $\mathbb{E}\left[1-\frac{{\binom{n-c}{k}}}{\binom{n}{k}}\right]$, which is the same with Pass@$k$ (Chen et al., 2021).
@@ 3.2 Results and Analysis
Figure 3: The EM@1 results for token-level code completion from VersiCode: (a1) Comparison with existing benchmark datasets, (a2) Performance grouped by data sources, and (b) Performance grouped by API lifecycle.
@@
Even token-level code completion is challenging. We present the EM@1 results for token-level code completion on VersiCode, sorted by release time (highlighted in green, see Figure 3-a1). Compared to the Pass@1 results on HumanEval (blue) and MBPP (orange), all models perform significantly worse on VersiCode (green). This result indicates the difficulty in disambiguating and recalling version-specific library usage. It is important to note that larger and more recent models, such as GPT-4o (M13) and LLaMA3-70B (M12), demonstrate significantly superior performance compared to other models. (See Appendix H for the error analysis of GPT-4o.) However, a substantial performance gap of at least 15 points remains when compared to HumanEval and MBPP (detailed in Appendix D.1). This indicates that state-of-the-art LLMs still struggle to deliver satisfactory results, even for the simplest token-level completion tasks.
@@
Differences in LLM performance across different data sources. Figure 3-a2 presents the EM@1 results for token-level code completion on VersiCode, categorized by data sources. Among the three data sources, most models perform significantly better on Stack Overflow, especially compared to handling source code from downstream applications. This discrepancy may be attributed to the greater diversity found in downstream applications, which demands a more robust capability to address varied challenges. This may also indicate that Stack Overflow is heavily represented in the pre-training data of LLMs, increasing the likelihood of data leakage. Similar to Figure 3-a1, GPT-4o (M13) and LLaMA3-70B (M12) stand out as outliers, excelling in handling downstream applications, which may increase the likelihood of models memorizing specific content. Full numeric results are provided in Appendix D.1.
@@
Challenges in casual intermediate library versions. We present the token-level EM@1 results for the token-level code completion task, categorized by lifespan features: addition (in blue), deprecation (in orange), and general (referring to intermediate versions; in green), as shown in Figure 3-b. Most models perform well in cases of addition and deprecation, likely because newly added or deprecated APIs are often emphasized in documentation and by the community. However, most models struggle with reasoning and adapting to intermediate versions. As shown in Figure 3-a2, models like LLaMA3-70B excel in downstream applications and handle intermediate versions more effectively, likely due to the diversity of use cases they encounter.
@@
Figure 4: The EM@1 performance for token-level code completion, grouped by year (2015-2023), with a histogram of data distribution for each year.